# Applications of Generative AI (GAI) for Mobile and Wireless Networking: A Survey

### v1

@@ II-A History of Generative AI
AI has a long history and has developed a lot over many years. Due to big investments from governments and large companies, AI has reached important milestones and is used in almost every industry, including infotainment, finance, advertising, defense, agriculture, and telecommunications.
In general, AI techniques and their applications are designed by researchers in the field of mobile networking. AI-driven algorithms are used by researchers to analyze and optimize data traffic, improving network efficiency and performance, while automating network management tasks like configuration and troubleshooting [45]. These algorithms also facilitate the dynamic allocation of network resources based on real-time demand and usage patterns, optimizing resource utilization and enhancing network capacity and scalability [46]. To ensure network security, AI-powered IDS analyzes network traffic patterns in real time, identifying and mitigating potential threat, and supporting network resilience against cyber attacks and unauthorized access [47]. Additionally, AI-based encryption and authentication mechanisms ensure data confidentiality and integrity, protecting sensitive information transmitted over mobile networks from unauthorized access and tampering [48].
@@
Ref. Characteristics Description Advantages Challenges
@@
Despite showing great potential in various domains, conventional AI approaches may not be effective in highly dynamic and complex environments (e.g., future mobile and wireless networking), as discussed in Section I. To overcome this drawback, the potential of GAI has recently been demonstrated by advances in VAEs [59] and GANs [60]. These innovations were based on Deep Neural Network (DNN) architectures capable of learning generative models for complex data types like images. In 2017, Google introduced the Transformer architecture [61] that further enhanced the capability of GAI in learning generative models. This led to the development of the first Generative Pre-trained Transformer (GPT), as known as GPT-1, in 2018. Since then, GAI has received attention from both academia and industry, with various advantaged techniques and applications, e.g., DALL-E and GPT-4. The advantages and challenges of GAI use cases can be summarized in Table II. In the following, we present the fundamentals of common GAI techniques together with their potential in mobile and wireless networking areas.
@@ II-B Primitive Structure of GAI
One of the core objectives of generative adversarial learning-enabled trust management services for Beyond Fifth-Generation (B5G)/Sixth-Generation (6G) wireless networks is to enhance intelligence and security while ensuring reliable and real-time network security and service quality optimization [15]. The core elements inherent in the basic structure of GAI models empower them to provide robust and reliable services for both modern wireless service providers and users.
@@ II-B1 Variational Autoencoders
Variational Autoencoders (VAEs), illustrated as in Fig. 3, play a crucial role in GAI by facilitating the generation of realistic and high-quality data. They are applied across diverse domains like Computer Vision (CV), NLP, and mobile networking. Especially in situations with limited labeled data, autoencoders are highly valuable as they can generate synthetic data samples for training purposes [62]. Furthermore, VAEs contribute significantly to dimensionality reduction and feature extraction, leading to a more compact and structured representation of complex data. Fundamentally, the encoder compresses data into a lower-dimensional representation, namely latent space, while the decoder module reconstructs the data from the latent space, restoring it to the original form. Once trained, the decoder becomes capable of generating new data points by sampling from the latent space [63]. This remarkable capability makes VAEs highly valuable for tasks such as data synthesis, text generation, anomaly detection, and wireless networks.
Figure 3: An illustration of an autoencoder-based GAI. Here, the generated data is guaranteed to be robust to noise as the hidden layers are being injected by noise during the training to simulate the channel noise of the wireless channel.
II-B2 Diffusion Models
An alternative approach to generative modeling is offered by diffusion models, which draw inspiration from non-equilibrium thermodynamics [64]. These models use a Markov chain of diffusion steps to gradually add random noise to the input data. Then, they learn to reverse this diffusion process to create desired data samples from the noise. Unlike VAE or flow models, diffusion models use a fixed training procedure and have high-dimensional latent variables that match the dimensionality of the original data. Table III highlights the advantages and limitations of GAI models in mobile networking applications.
GAI models are intended to generate new data samples that resemble a given training dataset. A GAN trained on a dataset of landscape images, for example, can produce new landscape images that resemble the training data but are not exact replicas of any particular image [65]. In GAI, there are several methods to produce data, such as likelihood-based models, energy-based models, and two-star models [17]. Estimating the probability distribution of the training data is the main goal of likelihood-based models, such as the Gaussian Mixture Model (GMM). GMM assumes that the data are produced by combining different Gaussian distributions. Data collection entails selecting samples from the learned distribution via a trained GMM. For instance, we can train a GMM with a handwritten number dataset; after learning its features, the data can be configured to produce novel samples [66].
Energy-based models define an energy function that assigns low energy to low-resolution samples from the training set and high energy to other samples. To find low-energy samples that are probably similar to the training data, one can adopt Markov Chain Monte Carlo (MCMC) techniques to search the sample space [67].
An alternative method involves incorporating GANs into diffusion models. In this approach, the generator endeavors to mislead the discriminator, while the discriminator strives to accurately classify the samples. To achieve this, these two models are trained in opposition to each other [68].
II-B3 Generative Adversarial Networks
@@
GANs [60] were introduced to generate realistic synthetic data. The key rationale behind GANs lies in their unique architecture, which consists of two neural networks: the generator and the discriminator, engaged in a competitive game. The discriminator distinguishes genuine and generated data, while the generator creates synthetic data samples that closely resemble real data. This interplay drives an adversarial training process where the generator aims to deceive the discriminator and enhance the quality of its output. GANs have diverse applications in generative AI, including text generation, data augmentation, and image synthesis. They excel in generating a wide range of samples and capturing complex data distributions. Furthermore, the incorporation of techniques such as conditional GANs [69, 70], style transfer [71, 72], and progressive training [73] enhance the control and manipulation of the generated outputs. These advancements have made GANs indispensable in various industries, including art, fashion, and entertainment. Such methods also facilitate the production of diverse and realistic data, where their adversarial training process can capture complex data distributions and adapt to different applications. The use of GANs has transformed the field of GAI, enabling the creation of unique and high-quality content across multiple domains.
@@ II-C Multi-modal GAI
Multi-modal GAI encompasses a range of tasks, including but not limited to image recognition, semantic segmentation, NLP, speech synthesis, and emotion detection. Through the integration of diverse modalities like visual, auditory, and textual information, multi-modal GAI is capable of capturing a holistic representation of the world, mirroring the way humans perceive and interpret multiple sources of information concurrently [74]. In what follows, we explain how GAI deals with multi-tasking problems.
@@ II-C1 Meta-learning GAI
Using meta-learning [75], GAI systems can achieve fast adaptation by leveraging the task-specific gradient trajectories. This approach enables GAI systems to acquire knowledge more efficiently and transfer it across diverse domains [76], while also possessing the capability to continue learning [77, 78]. Additionally, meta-learning addresses the challenge of continuous learning by mitigating the issue of catastrophic forgetting, allowing GAI to retain previously acquired knowledge while continuously acquiring new knowledge. For example, the authors in [79, 80] propose a data acquisition framework based on GANs and meta-learning to reduce the need for collecting a large number of real samples. This framework generates wireless data samples for DL models with the same performance as models trained with more real samples. The research in [81] proposes a different way of leveraging meta-learning into GAI design. Instead of applying meta-learning to quickly adapt the GAI model from different domains, MetaMask leverages meta-learning to efficiently adapt the joint GAI model via trajectory estimation from the GAI sub-model. Fig. 4 shows the taxonomy of meta-learning and its application using GAI.
@@ II-C2 Multi-task GAI
Multi-task GAI refers to the ability of an AI system to effectively perform and generalize across a wide range of tasks [82, 83, 84]. By leveraging the knowledge and skills gained from one task and applying them to related or similar tasks, multi-task GAI utilizes transfer learning to improve performance across various domains [85, 86]. Moreover, these systems extract shared representations and beneficial characteristics that contribute to improved performance on new and unexplored tasks. Collaborative learning from multiple tasks reduces training time and enhances computational efficiency, optimizing resource utilization. Another advantage of multi-task GAI is domain adaptation, enabling the system to adjust its knowledge and skills to different domains or tasks [87]. This adaptability empowers the system to achieve high performance in unfamiliar or dynamic environments. Additionally, multi-task GAI facilitates continuous learning by gradually introducing new tasks while minimizing the forgetting of previously mastered tasks. This dynamic approach allows ongoing knowledge upgrades and long-term proficiency across diverse disciplines.
@@ II-D Representative GAI Models
Across a diverse range of tasks and domains, several representative GAI models have demonstrated impressive capabilities. One such example is GPT-3, a renowned model developed by OpenAI. GPT-3 stands out for its exceptional language processing abilities, excelling in tasks like language generation, translation, summarization, and question-answering [88]. In addition, OpenAI’s DALL-E model is a specialized form of GAI that focuses on generating images from textual descriptions, effectively transforming natural language prompts into accurate visual representations [89]. It enables tasks such as image categorization, object detection, and text-based image synthesis. Contrastive Language-Image Pre-training, another creation from OpenAI, integrates language and image understanding in a cohesive manner [90]. These GAI models exemplify the potential for developing intelligent systems that excel across various activities and domains, showcasing cutting-edge progress in the field of AI.
III Application of Generative AI in Network Management
As modern networks exponentially grow in complexity and scale, traditional management approaches face challenges in ensuring efficiency, security, and adaptability. Integrating GAI intro mobile networks help overcome these circumvents, approaching the era of intelligent network management. To explain how GAI works, this discusses its particular applications from network management, network routing, and resource allocation to channel estimation. Then, we conclude this section with useful learned and key takeaways.
III-A Network Management
@@
Software-Defined Networking (SDN) represents a pivotal convergence point between GAI and network management. To be specific, SDN’s separation of the control plane and data plane facilitates centralized network management and programmability. Integrating GAI into SDN frameworks empowers network administrators with enhanced automation, intelligent resource allocation, and adaptable network optimization. Specifically, GAI is in charge of real-time decision-making for effective traffic routing, load balancing, and Quality-of-Service (QoS) management, enabling SDN controllers to dynamically learn from and adapt to changing network conditions. For example, DGMs are used in [24] to address the issues of data scarcity, limited flexibility, and dynamic network states using a diffusion model to generate effective contracts. In [91], NetGPT is proposed to employ LLMs at the edge and the cloud in conjunction with location-based information to speed up personalized procedures and cloud interaction, enhancing intelligent network management and orchestration.
@@
Network slicing, a vital aspect of network management, has great potential to leverage GAI and opens a new chance to customize network instances within a shared infrastructure [92]. On this basis, GAI can exploit this property to perform network data analysis, pattern prediction, and generation of optimized network slices tailored to specific application requirements. In this way, mobile networks can dynamically allocate resources, balance workloads, and adapt to changing traffic demands, contributing to better user experiences, increased network efficiency, and greater flexibility in service provisioning [93].
@@ III-B Resource Allocation
In the context of dynamic resource allocation, GAI contributes to dynamically changing network conditions in real time through continuous monitoring and analysis of network data. This enables efficient utilization of resources, ensuring optimal performance and responsiveness to the evolving needs of various network factors such as network load, user demands, and QoS demands. It also helps systems use proper resource allocation strategies to ensure an optimal allocation that aligns with network objectives and user expectations.
@@
III-C Network Routing
To deal with the dynamics and uncertainty of network conditions, introducing GAI into network routing can be an innovative alternative to conventional AI approaches. Specifically, GAI models can assess the efficacy of various routing algorithms and optimize routing choices for certain network objectives by simulating and creating synthetic network scenarios. This guarantees effective resource utilization, low latency, and better overall network performance. Moreover, by analyzing network data to spot potential weaknesses and learn from historical security incidents, GAI helps navigate and enhance routing patterns to reduce the possibility of assaults or unauthorized access, proactively mitigating security concerns.
A typical example of this feature is the adoption of DGMs in network economics to improve the efficiency of wireless network management by exemplifying and generating effective contracts for incentivizing mobile AIGC services [95]. Another example of GAI in routing solutions for different network status distributions and topology is exploiting a transfer RL algorithm to improve training efficiency by rapidly transferring knowledge [96].
III-D Channel Estimation

### v2

@@ II-A History of Generative AI
Generative AI is rooted in ML and DL algorithms, which have evolved since the 1950s. The term “machine learning” was coined by Arthur Samuel in 1952 when he developed an algorithm for playing checkers. In 1957, Frank Rosenblatt, a psychologist at Cornell University, invented the first “neural network” capable of being trained, known as the Perceptron, which had a single layer with adjustable thresholds and weights. A decade later, in 1961, Joseph Weizenbaum created ELIZA, an early prototype of GAI and one of the first chatbots. ELIZA was designed to simulate empathetic conversation using natural language processing. During the 1960s and ’70s, Ann B. Lesk, Leon D. Harmon, and A. J. Goldstein made significant strides in computer vision (CV) and pattern recognition by identifying 21 specific human signs. In the 1970s, Seppo Linnainmaa introduced the term “backpropagation”, a technique involving three main steps: processing at the output end, distributing backward, and moving through the network’s layers for training and learning. This method laid the groundwork for training Deep Neural Networks (DNNs).
The distinction between ML and AI became evident during two “AI winter” periods. The first AI winter, from roughly 1973 to 1979, saw little progress despite high expectations. However, ML continued to advance, proving cost-effective and useful for business applications like automated phone systems. In 1979, Kunihiko Fukushima developed the first multilayered Artificial Neural Network (ANN), enabling computers to recognize visual patterns and handwritten characters. The second AI winter, from around 1984 to 1990, was marked by slow development and widespread skepticism, leading to reduced funding for AI and DL research. In the late 1980s, the integration of Metal Oxide Semiconductors and Very Large Scale Integration technology resulted in efficient ANNs. In 1989, Yann LeCun’s neural networks for recognizing handwritten ZIP codes marked a significant breakthrough. This progress continued with the presence of long short-term memory (LSTM) networks in 1997 and advanced GPUs by Nvidia in 1999. These advancements paved the way for launching the first digital virtual assistant, Siri, on October 4, 2011.
The use of chatbots spiked, along with the introduction of VAEs in 2013 [51] and GANs in 2014 [52], both of which were built on top of DNN architectures capable of learning generative models for complex data types such as images. Upon this progress, in 2017, Google introduced the “Transformer” architecture [53] to enhance the robustness of GAI in learning generative models, followed by the first Generative Pre-trained Transformer (GPT), as known as GPT-1, in 2018. Four years later, OpenAI introduced ChatGPT, a GAI model leveraging LLM, achieving new levels of performance in research, writing, and content creation. Since then, the use of GAI has spread from academia to industry, with diverse applications including infotainment, finance, advertising, defence, agriculture, and telecommunications. However, everything has two sides, and so does GAI. Table II covers the advantages and challenges of GAI use cases. To better understand the core value of GAI, we next go into the fundamentals of GAIs.
@@
Ref. Attributes Description Advantages Challenges
@@ II-B Primitive Structure of GAI
The evolution of GAI spans a long history as mentioned earlier, and its recent variants have also become diverse for different applications. However, a basic architectural block of GAI includes three fundamental components: VAE, diffusion model, and GAN. The advantages and limitations of VAEs, diffusion models, and GANs are summarized in Table III.
@@ II-B1 Variational Autoencoders
VAE, a family of probabilistic graphical models and variational Bayesian methods, is an ANN introduced by Diederik P. Kingma and Max Welling [51], as shown in Fig. 3. In VAEs’ architecture, the encoder compresses data into a lower-dimensional representation, namely latent space. At the same time, the decoder module reconstructs the data from the latent space, restoring it to its original form. Once trained, the decoder can generate new data points by sampling from the latent space [64]. In GAI frameworks, VAEs are in charge of generating realistic and high-quality data, making them extremely valuable in situations with limited labeled data due to their ability to crease synthetic data samples for training purposes [65]. Besides, VAEs contribute significantly to dimensionality reduction and feature extraction, leading to a more compact and structured representation of complex data.
Figure 3: Summary of GAI-involved information: The history of GAI, its primitive structure, variation of GAI, and representative GAI models.
II-B2 Generative Adversarial Networks
@@
GANs were introduced to generate realistic synthetic data [52]. The key rationale behind GANs lies in their unique architecture, which consists of two neural networks: the generator and the discriminator, engaged in a competitive game. The discriminator distinguishes genuine and generated data, while the generator creates synthetic data samples that closely resemble real data. This interplay drives an adversarial training process where the generator aims to deceive the discriminator and enhance the quality of its output. As a result, GANs can produce diverse and realistic data since their adversarial training process can capture complex data distributions and adapt to different applications. Especially, GANs also demonstrate their unique difference from VAEs and diffusion models by their enhanced ability to control and manipulate the generated outputs through techniques such as conditional GANs [66, 67], style transfer [68, 69], and progressive training [70].
II-B3 Diffusion Models
Diffusion models are a class of probabilistic generative models, which draw inspiration from non-equilibrium thermodynamics [71]. Technically, diffusion models consist of two distinct stages. In the first phase, a series of diffusion steps is applied to incrementally introduce random noise into the input data, where the noise magnitude varies at each step. In the subsequent phase, this process is reversed: the model iteratively learns to remove the noise, progressively reconstructing the data samples from the noisy input. As a result, during inference, data are generated by gradually reconstructing them, beginning from an initial state of random white noise. Diffusion models can be classified into three primary subcategories, with the main differences lying in the way they are used to progressively denoise the data.
Score-based Generative Model (SGM)
SGM is proposed by [72] to address the challenges of generative models (e.g., VAEs, GANs), including: 1) mode collapse, where the model fails to capture the full diversity of data distribution, 2) instability due to the adversarial training procedure, and 3) GAN objective function is not suitable for evaluating and comparing among baselines. Besides, SGM also guides a new principle for generative modeling based on estimating and sampling from the (Stein) score [73] of the gradient of the log-density function at the input data point, i.e., vector field pointing in the direction where the log data density grows the most. The rationale behind this approach is to leverage the Langevin dynamics to move from a random sample $x_{0}$ to samples $x_{N}$ in regions of high density $p(x)$ [72].
Denoising Diffusion Probabilistic Models (DDPM)
Neural network $\theta$ approximates the diffusion steps by leveraging $p_{\theta}(x_{t-1}|x_{t})=\mathcal{N}(x_{t-1};\mu_{\theta}(x_{t},t),\Sigma_{\theta}(x_{t},t))$ to progressively denoise the data $x_{t}$ at time step $t$, with a focus on learning to predict the noise trajectory with mean $\mu_{\theta}(x_{t},t)$ and covariance $\Sigma_{\theta}(x_{t},t))$. In ideal cases, the model $\theta$ can be trained with a maximum likelihood objective so that the model inference $p_{\theta}(x_{0})$ is as identical to each training example $x_{0}$. However, having $p_{\theta}(x_{0})$ is quite intricate due to the required marginalization of all possible reverse trajectories. This promotes the study on leveraging VAE to minimize the variational lower-bound of the negative log-likelihood instead [71], fixing $\Sigma_{\theta}(x_{t},t))$ to a constant value combined with rewriting $\mu_{\theta}(x_{t},t)$ as a function of noise [74], and reformulating the diffusion process as a non-Markovian process [75]. Compared to [71] and [74], the DDIM model in [75] achieves faster data generation with minimal compromise in data quality, for its ability to express the denoising procedure as an implicit function instead of a stochastic process.
Stochastic Differential Equations
Similar to the previous two methods, the approach presented in [76] gradually transforms the data distribution $p(x_{0})$ into noise. However, it extends the previous methods by considering the diffusion process as continuous, which is modeled as the solution to a Stochastic Differential Equation (SDE). In particular, [76] uses a neural network $\theta$ to estimate the score functions, similar to [72], and generates samples from $p(x_{0})$ using SDE solvers.
@@ II-C Multi-modal GAI
Multi-modal GAI is an advanced level of GAI that allows one to handle various tasks with different requirements. This is achieved by integrating diverse modalities like visual, auditory, and textual information [77]. However, due to the different uses of real-world scenarios, multimodal GAI can be differentiated into meta-learning GAI and multi-task GAI. The following discussion will explore how each method addresses specific scenarios with its methods and use cases.
@@ II-C1 Meta-learning GAI
Meta-learning GAIs refer to “learning to learn” models that aim to quickly adapt to new tasks with minimal data constraints by leveraging prior knowledge. This is achieved by training a model on various combinations of tasks (e.g., text, image, and audio) to learn a meta-knowledge or strategy that can be applied to new or unseen tasks. This feature becomes particularly useful in few-short or zero-short learning scenarios.
For example, GAI systems with meta-learning can achieve fast adaptation by leveraging the task-specific gradient trajectories [78]. This approach enables GAI systems to acquire knowledge more efficiently and transfer it across diverse domains [79], while also possessing the capability to continue learning [80, 81]. Additionally, meta-learning also addresses the challenge of continuous learning by mitigating the issue of catastrophic forgetting, allowing GAI to retain previously acquired knowledge while continuously acquiring new knowledge or reducing the need for collecting a large number of real samples [82, 83]. Especially, instead of applying meta-learning to quickly adapt the GAI model from different domains, a different way of leveraging meta-learning, called MetaMask, is proposed in [84], with the focus of leveraging meta-learning to efficiently adapt the joint GAI model via trajectory estimation from the GAI sub-model. The taxonomy of meta-learning GAIs can be summarised by [85].
@@ II-C2 Multi-task GAI
Multi-task GAI refers to the ability of an AI system to effectively perform and generalize across a wide range of tasks [86, 87, 88]. It utilizes transfer learning to improve performance across various domains by leveraging the knowledge and skills gained from one task and applying them to related or similar tasks [89, 90]. Moreover, it can extract shared representations and beneficial characteristics to improve performance on new and unexplored tasks. Collaborative learning from multiple tasks reduces training time and enhances computational efficiency, optimizing resource utilization. Another advantage of multi-task GAI is domain adaptation, enabling the system to adjust its knowledge and skills to different domains or tasks [91]. This adaptability empowers the system to achieve high performance in unfamiliar or dynamic environments. Additionally, multi-task GAI facilitates continuous learning by gradually introducing new tasks while minimizing the forgetting of previously mastered tasks. This dynamic approach allows ongoing knowledge upgrades and long-term proficiency across diverse disciplines.
@@ II-D Representative GAI Models
Across a diverse range of tasks and domains, several representative GAI models have demonstrated impressive capabilities. For example, GPT-3, a renowned model developed by OpenAI, stands out for its exceptional language processing abilities, excelling in tasks like language generation, translation, summarization, and question-answering [92]. OpenAI’s DALL-E model is a specialized form of GAI that focuses on generating images from textual descriptions, effectively transforming natural language prompts into accurate visual representations [93]. It enables tasks such as image categorization, object detection, and text-based image synthesis. However, GPT-3 and DALL-E require overparameterization, making them unsuitable for deployment in IoT systems.
Inspired by this, Contrastive Language-Image Pre-training (CLIP) [94] and Bootstrapping Language-Image Pre-training (BLIP) [95] have emerged as promising candidates for the future of GAI in IoT networks due to twofold: 1) Lightweight architectures, and 2) Pretrained models, allowing their parameters to be effectively transferred to a variety of applications such as conditional text to image (ControlNet) [96], Zero-shot image to image (Pix2Pix-Zero) [97]. These GAI models exemplify the potential for developing intelligent systems that excel across various activities and domains, showcasing cutting-edge progress in the field of AI.
Figure 5: Illustration of GAI applications in network management, including network control, resource allocation, and channel estimation.
III Application of GAI in Network Management
This section reviews the applications of GAI in network management, including network control, resource allocation, and channel estimation. Fig. 5 illustrates the applications of GAI in network management. Then, we conclude this section with lessons learned and key takeaways.
III-A Network Control
In the context of network softwarization, network control can be classified into three levels: Software-Defined Networking (SDN), network slicing, and network routing [98]. SDN centralizes and automates network control, allowing for dynamic adjustments and efficient resource management. Network slicing leverages SDN to create customized, isolated virtual networks tailored to specific needs, enhancing flexibility and efficiency. The centralized management of SDN facilitates network routing, a fundamental component, which ensures the best path selection for data transmissions.
Software-Defined Networking
@@ III-A Network Control
SDN separates the network’s control plane, which makes decisions about traffic routing, from the data plane, which forwards traffic to the selected destination. Particularly, SDN introduces dominant features that enable billions of IoT-connected devices to be controlled and accessed remotely [99]. Thus, integrating GAI into SDN frameworks empowers network administrators with enhanced automation, intelligent resource allocation, and adaptable network optimization, thereby meeting the IoT requirements in an efficient, scalable, seamless, and cost-effective manner. More specifically, GAI is in charge of real-time decision-making for effective traffic routing, load balancing, and Quality-of-Service (QoS) management, enabling SDN controllers to learn from and adapt to changing network conditions dynamically. For example, using a diffusion model to generate effective contracts in [26] can address the issues of data scarcity, limited flexibility, and dynamic network states. In [100], NetGPT is proposed to employ LLMs at the edge and the cloud in conjunction with location-based information to speed up personalized procedures and cloud interaction, enhancing intelligent network management and orchestration.
Network Slicing
@@
Network slicing is used primarily in Fifth-Generation (5G) networks and Sixth-Generation (6G) to create multiple virtual networks on a shared physical infrastructure. Especially, in the IoT field [101], network slicing can accommodate different network needs of heterogeneous IoT applications through dedicated slicing. This presents a promising research direction, where leveraging GAI helps to enhance customizing network instances within shared IoT infrastructures [102]. For example, GAI can exploit the virtual properties to perform network data analysis, pattern prediction, and generate optimized network slices tailored to specific requirements. In this way, wirelessly intelligent networks can dynamically allocate resources, balance workloads, and adapt to changing traffic demands, contributing to better user experiences, increased network efficiency, and greater flexibility in service provisioning [103].
Network Routing
Network routing typically consumes significant bandwidth and resources, particularly when selecting optimal paths to ensure efficient data transmissions from the source to the destination in large-scale IoT environments, where frequent updates are necessary. Introducing GAI into network routing can enhance the efficacy of choosing and optimizing routing algorithms for certain network objectives by simulating, creating, and analyzing synthetic network scenarios. An example of this is exploiting a one-shot conditional generative routing model to perform one-shot routing to the pins within each network, and the order in which the networks need to be routed is learned adaptively [104]. Another example of GAI in routing solutions for different network status distributions and topology is exploiting a transfer RL algorithm to improve training efficiency by rapidly transferring knowledge [105].
@@ III-B Resource Allocation
Resource allocation involves distributing available resources such as bandwidth, power, and time slots to optimize network performance for various users and services. In the context of dynamic wireless IoT networks where access entities frequently arrive and depart in short periods and changes in wireless mediums are unpredictable, GAI contributes to dynamically changing network resources in real-time through continuous monitoring and analysis of network traffic. This ensures optimal performance and responsiveness to the evolving needs of various network factors such as network load, user demands, and QoS demands. It also helps systems use proper resource allocation strategies to ensure an optimal allocation that aligns with network objectives and user expectations.
@@
III-C Channel Estimation